OpenStudy (princesssleelee):

Which graph represents the function of f(x) = the quantity of 9 x squared minus 36, all over 3 x plus 6? graph of 3 x plus 6, with discontinuity at 2, 12 graph of 3 x plus 6, with discontinuity at negative 2, 0 graph of 3 x minus 6, with discontinuity at negative 2, negative 12 graph of 3 x minus 6, with discontinuity at 2, 0

OpenStudy (aihberkhan):

Okay. Can you show me your work or at least try to solve? It is exactly like the previous question that I helped you with, which took a lot of explaining :) @princesssleelee

OpenStudy (princesssleelee):

well first find the common factor in the denominator.

OpenStudy (princesssleelee):

which is 3. x+2

OpenStudy (princesssleelee):

correct?

OpenStudy (aihberkhan):

Right... \(3(x + 2)\)

OpenStudy (aihberkhan):

So we are simply factoring again. We would take \(a^2−b^2=(a+b)(a−b)\) Now lets look at the numerator, shall we? \(9x^2−36\) We know that \(9 = 3^2\) and that \(3^2 x^2 = (3x)^2\) Then, \(36 = 6^2\) Now, when looking at this we do have a difference of squares, correct? Which is... \((3x)^2 - 6^2\)

OpenStudy (princesssleelee):

yes

OpenStudy (aihberkhan):

So now lets just factor and substitute in our numbers. \( a^2−b^2=(a+b)(a−b)\) We know that \(a = (3x)\) and \(b = 6\) Right?

OpenStudy (princesssleelee):

right

OpenStudy (aihberkhan):

Okay, now lets just substitute in our values. \((3x)^2−6^2=((3x)+6)((3x)−6)\) Now let's "explore" and "experiment" a bit. If we were to rewrite the numerator and change it to \((3x + 6)(3x - 6)\) and look at the fact that the denomination has a factor of \(3x + 6\), we can cancel those out... right?

OpenStudy (princesssleelee):

right

OpenStudy (princesssleelee):

because the signs are opposite meaning they cancel.

OpenStudy (aihberkhan):

Good Job! Now let's complete the "experiment" shall we? So if we were to substitute that part in, it would look like this: \(f(x)=(3x+6)(3x−6)3x+6\) We can simply cancel these two factors out. However, \(3x + 6\) can't equal 0. We know that \(3x + 6 = 0\) when \(x = -2\). This means that we will not have any point at \(x = -2\) Now we can cancel the factors to equal 1. \(f(x)=3x−6x≠−2\) Now can you try finding the graph of those answer choices?

OpenStudy (princesssleelee):

i know it is not A

OpenStudy (aihberkhan):

Okay. Continue...

OpenStudy (princesssleelee):

its below -2 so C and D are my options, right?

OpenStudy (aihberkhan):

Right! Now what...

OpenStudy (princesssleelee):

well, i have to determine the plotted point

OpenStudy (aihberkhan):

You see... You want the one where the point is on -2. Can you figure it out now?

OpenStudy (princesssleelee):

leading to graph C! its x-axis is on -2

OpenStudy (aihberkhan):

Awesome!

OpenStudy (aihberkhan):

Hope this helped! Have a great day! :) A medal would also be appreciated! Just click best response next to my name! Also, a fan would be appreciated as well! Just hover over my icon and click "Become A Fan". This will allow you to see every time I am online! :) If you see that I am online and need help with a question, just tag me in your question! @princesssleelee

OpenStudy (princesssleelee):

Already a fan! and i truly appreciate it !!!

OpenStudy (aihberkhan):

No problem! If we are finished... please close this post :)

OpenStudy (princesssleelee):

the last one we did was not letter D....

OpenStudy (tkhunny):

Great. Why don't you prove what it WAS and how you managed the appropriate response.

OpenStudy (aihberkhan):

This one?

OpenStudy (princesssleelee):

Yes

OpenStudy (aihberkhan):

We didn't choose letter D for this one..

OpenStudy (princesssleelee):

oh goodnesss.. thats the one i selected :(

OpenStudy (aihberkhan):

Aw... we chose C! :(

OpenStudy (princesssleelee):

OH NO the one before

OpenStudy (princesssleelee):

the last post

OpenStudy (princesssleelee):

c was correct :)

OpenStudy (aihberkhan):

Awesome!

OpenStudy (princesssleelee):

its okay, im not going to retake it.. that took way too long!

OpenStudy (princesssleelee):

an 80% will jut have to do, even though the rest are all 100%'s lol ;)